OpenStudy (anonymous):

solve each system of equations by the substitution method. Tell whether each system is inconsistent or consistent. If the system is consistent, tell whether it is independent or dependent 5x-4y=11 3x-y=1

OpenStudy (anonymous):

5x-4y=11 ... (1) 3x-y=1 ....(2) take one of the equations and solve for a variable in place of another variable: eg. take equation (2) for example 3x -y =1 solve for x add y to both sides => 3x -y +y = 1 +y => 3x =1+y => x = (1+y)/3 ..... (3) Next substitute for the new x value into equation (1) thus equation (1) becomes: 5x -4y = 11 => 5 * (1+y)/3-4y = 11 => 5 +5y -12y = 33 ( after multiplying by 3 on both sides) => -7y = 28 => y = -4 Now substitute the new value of y in equation (3) above Thus x = (1+y)/3 ..(3) =>x =(1-4)/3 = -1 Thus the solution is: x = -1 y = -4

OpenStudy (anonymous):

The system is consistent and independent